Is it unethical to purposely distort communication to get favourable outcome?
The answer depends on the situation.
In baseball, when the catcher signals the pitcher with finger signals, he is concealing his communication from the other team.
That is not unethical.
Without such justification, the answer seems to be yes.
What is the sample space when selecting a prime number between 10and30?
{11,13,17,19,23,29} are all the prime numbers between 10 and 30
When a two dice is rolled what is the probability the two number will have a sum of 8?
5/36
Out of the 36 possible and equally-probable outcomes (1st die's number and 2nd die's number), 5 combinations give you a sum equal 8:
* 1st die gives a 2 and 2nd die gives a 6
* 1st die gives a 6 and 2nd die gives a 2
* 1st die gives a 3 and 2nd die gives a 5
* 1st die gives a 5 and 2nd die gives a 3
* 1st die gives a 4 and 2nd die gives a 4
Therefore the probability that the sum is greater than 10 equals 5/36.
What is identical fraternal twins?
Too bad. I don't know. Go to a more reliable source. I just proved that answers here can be half-assed. Now go cry in a corner. Yeah, yeah.
What does equally probable mean?
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
That there is the same chance of something happening as there is of something else happening. When tossing a coin, there is the same chance it will land on heads as there is that it will land on tails. So landing on heads and landing on tails are equally probable.
You divide the area of the shaded region by the area of the full circle. For example, if the radius of the shaded region is 2 meters, the probability would be 4pi / 36pi, or 1/9. If the shaded region is a 'slice' of the circle, the chance is just the fraction of the circle which the 'slice' is.
Why can coin flips be used to represent the possible genders of children in families?
They are a crude way to represent children's genders which depend on the assumptions that the probabilities are a half for either gender and that the probabilities are independent from one child to another.
In fact, neither assumption is true.
According to recent data, the probabilities of a child's gender is not 0.5 but males are more likely: prob = 0.5169.
Children's genders depend on parental ages and genes and so they are not independent events.
What is the probability of getting a factor of 5 if you roll 1 dice twice?
Assuming that it is a fair die, the answer is 5/9.
What is the Probability density function of Poisson distribution?
If a random variable X has a Poisson distribution with parameter l, then the probability that X takes the value x is
Pr(X = x) = lx*e-l/x! for x = 0, 1, 2, 3, ...
How does the standard deviation impact the margin of error?
another way to look at this could be to compare a study done on a particular medicine, when study is on a small group of participants versus a large group the results must be shown as proportional and can only be relevant as they are calculated as how many are made well to how many do not receive any favorable results. Many times when a study is done on the effects of a substance on another it must be shown that the margin of error is proportionate to the study. As for instance the drug which conquered polio was simply a vaccine made of the polio virus and had great affects on the disease. Although there was a margin of error that the US government felt was acceptable to allow the drug to be forcibly injected in all children of the late 50's.
When two dice are rolled the probability of getting a sum between 5 and 11?
With 2 six-sided dice, there are 36 outcomes. The answer will depend on the meaning of between. First, here's a table of the number of possibilities to obtain each sum:
So if between means 6 through 10, then add up all of the above and you have 13 outcomes which are not between 5 & 11. So there are (36-13=23) outcomes which are between.
So the probability is 23/36. If you meant to include so that it was 5 through 11, then add 6 outcomes, so it is 29/36.
What is the probability of drawing 4 red cards out of a deck of playing cards?
26/52 x 25/51 x 24/50 x 23/49 = 0.055
Am I doing someones homework?
What is mean by -3standard deviation in bmi?
BMI varies from person to person and one measure of this variability is the standard deviation.
Assuming the BMI is approximately normally distributed, only around 0.135% of the people will have results that are -3 sd or lower.
What is event whose probability is 1 because it happens every time?
The probability that a person will die is 1.
The probability is 1 that you will pick a club if all the spades, hearts and diamonds are removed from a 52-card deck.
What are the differences between probability and coincidence?
The difference between probability and coincidence are that probability means the odds or how likely some thing will happen or not. While coincidence means something that did happen already, and the likely hood of what caused this to happen either being rare or common.
What is the probability of your birthday falling on a weekend?
The answer depends on where in the world you are. In some places, where hospital [or maternity unit] births are more popular, the medical staff prefer to deliver babies during weekdays so that they can have time with their own families at weekends. As a result, the distribution of births is not uniform across days of the week.
What is the probability of not rolling a sum of 2 4 or 6 with two fair dice?
There are 36 possible outcomes when throwing 2 dices : 6 for the first dice (1,2,3,...6) and 6 for the second dice
Let say that in the sums below the first number is the result of the first dice and the second number is for the result of the second dice:
2 = 1 + 1: 1 possibility
4 = 2 + 2 = 1 + 3 = 3 + 1: 3 possibilities
6 = 1 + 5 = 5 + 1 = 2 + 4 = 4 + 2 = 3 +3: 5 possibilities
1 + 3 + 5 = 9 possibilities to get 2 or 4 or 6
So the probability to get a sum of 2 or 4 or 6 is 9/36=1/4
And the probabilty of NOT rolling a sum of 2 or 4 or 6 is 1-1/4=3/4
Slightly under 50% - (.4925, to be exact).
----------------------------------------------------------------------------------------------------
Another opinion.
The probability that 3 balls are red is;
P(3R,2W) =
(5/11)(4/10)(3/9)(6/8)(5/7)[5!/(3!2!)] =
0.324675...
The probability that 4 balls are red is;
P(4R,1W) =
(5/11)(4/10)(3/9)(2/8)(6/7)[5!/(4!1!)] =
0.064935...
The probability that 5 balls are red is;
P(5R) =
5!6!/11! =
0.0021645...
The probability that at least 3 will be red balls is;
P(at
least 3R) =
P(3R,2W) + P(4R,1W) + P(5R) ~ 0.3918 ~ 39.2%
There are 50,400 ways and the probability (or probibility, as you prefer) of that answer being correct is 1.
What is the probability that when a fair 6 sided die is rolled a prime number faces up?
On a single roll, the probability is 1/2.
It is strangely worded like that, but the answer is yes.
What is Probability of Flipping a coin five times in a row and having it land on heads?
The probability of getting a heads on the first flip is 1/2.
Similarly, the probability on each subsequent flip is 1/2, since they are independent events.
The probability of several independent events happening together is the product of their individual probabilities.
How do find the probability of A complement given B complement?
With the information that is available from the question, it is impossible.
the colorblindness is usually not activited in a female body but is usually seen in male